OpenStudy (anonymous):

I hate to put questions on the wrong division, but I have a quiz tomorrow and no one has answered this in the chemistry department for a few hours. Stoichiometry 1 CuSO4 + 2 K --> 1 K2SO4 + 1 Cu How many moles of potassium are required to produce 4.15 moles of copper? How many grams of copper (II) sulfate are needed to react with 4.0 x 10^24 atoms of potassium?

OpenStudy (inkyvoyd):

ok, the potassium to copper ratio is 2:1 in terms of moles. so 2:1=x:4.15

OpenStudy (inkyvoyd):

does my reasoning (or lack there of) make sense?

OpenStudy (anonymous):

I guess I'm just confused as to how the ratio is determined.

OpenStudy (inkyvoyd):

the ratio of moles in a chemical equation is always constant. Moles have nothing to do directly with masses, but a mole is a quantity in units of an item. If you have a mole of potassium, you have 6.02*10^23 atoms of potassium. If you have a mole of K2SO4, than you have 6.02*10^23 units of K2SO4. Because chemicals react in units, you can use the ratio of a chemical equation to determine the amount of moles of one molecule/ionic bonding if you know the amount of moles of another.

OpenStudy (anonymous):

alright

OpenStudy (inkyvoyd):

how much of that made any sense?

OpenStudy (anonymous):

all of it minus the last sentance

OpenStudy (anonymous):

to produce 1 mole of Cu moles of K needed =2 To produce 4.15 moles of Cu moles of K needed=2*4.15=8.30

OpenStudy (anonymous):

that makes sense surjithayer

OpenStudy (inkyvoyd):

ok, so lets say you burn 1 mole of hydrogen gas. when you burn hydrogen gas you get water (since hydrogen plus oxygen=water). To figure out how much of it though you need the BALANCED chemical equation. I'll give it to you down here. 2H2+O2=2H2O in other words if you burn 2 moles (1.204*10^24 molecules) of hydrogen gas you need half the amount of moles, 1 mole (6.02*10^23 molecules) of oxygen gas. You will get 2 moles (1.204*10^24 molecules) of water.

OpenStudy (inkyvoyd):

does that make sense?

OpenStudy (anonymous):

a bit

OpenStudy (inkyvoyd):

what doesn't make sense?

OpenStudy (anonymous):

how you need half the amount of moles

OpenStudy (inkyvoyd):

let me start over. so let's say you have the reaction 2H2+O2=2H2O if I have 2 dozen hydrogen gas molecules than I need 1 dozen oxygen gas molecules to make 2 dozen water molecules. does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (inkyvoyd):

now, replace "dozen" with moles. A mole is the same kind of thing as a dozen, just a dozen refers to 12, while a mole refers to avogrado's number, 6.02*10^23

OpenStudy (anonymous):

okay